In organic chemistry, cis/trans isomerism describe
Answer #1

lnK_1 = lnA - \frac{Ea_1}{RT}

lnK_2 = lnA - \frac{Ea_2}{RT}

subtracting equation 1 from equation 2 we get

ln K_2 - lnK_1 = \frac{1}{RT}(Ea_1 - Ea_2)

ln \frac{K_2}{K_1} = \frac{1}{RT}(Ea_1 - Ea_2)

substituting the values we get

ln \frac{K_2}{K_1} = \frac{1}{8.314 * 353}(65000 - 27500)

ln \frac{K_2}{K_1} = 12.7775

\frac{K_2}{K_1} = e^{12.7775} = 354165.35 = 3.5 * 10^{5}

Hence the correct answer is Option A